OpenStudy (anonymous):

a music store is selling 2 different CDs by a local group. The group's first CD, awakening, sells for $16. their second CD, Encore, sells for $12. The stroe sells 240 of the group's CD's for a total of $3120. a. If "x" is the number of awakening CDs that were sold, write and expression for the number of encore CDs that were sold. b. write and solve an equation you can use to solve to find the number of Awakening and encore CDs sold. c. If the Awakening CDs were $18 instead of 16, would it still be possible for 240 CDs to be sold for a total of #3120? if so, how many of each were sold? justify

OpenStudy (anonymous):

do it yourself feather.

OpenStudy (anonymous):

BOOOOOOOOOO YOU. feather

OpenStudy (anonymous):

feather?

OpenStudy (anonymous):

LOL GENA . feather feather feather

OpenStudy (anonymous):

flutter me in the retriceole.

OpenStudy (anonymous):

a. y = 240 - x b. 16x + 12(240 - x) = 3120

OpenStudy (anonymous):

what about part c?

OpenStudy (anonymous):

b is not done yet. you may write a matrix: [1 1 240] [16 12 3120] in reduced row echelon form to find x and y you should get 60 awakening CDs and 180 encore CDs

OpenStudy (anonymous):

for c just find the reduced row echelon form of: [1 1 240] [18 12 3120] to get 40 awakening CDs and 200 encore CDs

OpenStudy (anonymous):

were in algebra two, so im not sure what echelon means, would you mind telling us in another way.

OpenStudy (anonymous):

adg - here is the answer to what you asked before: http://tinyurl.com/2uocfqh

OpenStudy (anonymous):

for b, just solve for the system of linear equations: x + y = 240 16x + 12y = 3120 for c, same thing: x + y = 240 18x + 12y = 3120

OpenStudy (anonymous):

thank you :)